OpenStudy (zmudz):

Prove that \(\sqrt{ \frac{2x^2 - 2x + 1}{2} } \geq \frac{1}{x + \frac{1}{x}}\) for \(0 < x < 1.\)

OpenStudy (thomas5267):

Square both side first?

OpenStudy (anonymous):

Fix up the RHS first.

OpenStudy (anonymous):

Note that \(x+1/x \geq 2\) (by arithmetic mean > geometric mean or simply by completing the square) Also, \(2x^2 - 2x + 1 = (1-x)^2 + x^2\) Now, use the fact that the root-means square is greater than the arithmetic mean, \(\sqrt{\frac{2x^2 - 2x + 1}{2}} = \sqrt{\frac{(1-x)^2 + x^2}{2}} \geq \frac{(1-x) + x}{2} = \frac{1}{2}\) Multiply the above two inequalities to get what is required.

OpenStudy (loser66):

@adxpoi I would like to know about the so-called "arithmetic mean> geometric mean". Would you please explain me?

OpenStudy (loser66):

The problem is easily proving by "square both sides" and do backward of the flow. I would like to learn a new method.

OpenStudy (anonymous):

Arithmetic mean is\[\frac{ x_1+x_2+x_3+...+x_n }{ n }\]while geometric mean is\[\sqrt[n]{x_1x_2x_3...x_n}\]

OpenStudy (anonymous):

Sure. The AM > GM inequality states that for \(n\) non-negative real numbers \(x_1, x_2,..., x_n\) the following holds: \[\frac{x_1+x_2+...+x_n}{n} \geq (x_1.x_2....x_n)^{\frac{1}{n}}\]

OpenStudy (loser66):

Thank you. I will read it carefully. @thomas5267

OpenStudy (anonymous):

A possible proof is by induction, first prove it for two numbers, easily done. Then try to extend it by induction.

OpenStudy (loser66):

@adxpoi Yes, I will. Thank you.

ganeshie8 (ganeshie8):

Alternatively you may show that the max value of right hand side is \(\frac{1}{2}\) and min value of left hand side is \(\frac{1}{2}\). They occur at different places, so the inequality follows.

ganeshie8 (ganeshie8):

thats exactly same as your first proof, but I was looking at it more from vertex of quadratic equation side.. :)

OpenStudy (loser66):

@ganeshie8 I looooooove your max/min method.!! Wow!!!

OpenStudy (anonymous):

Cumbersome, but doesn't require knowledge of the theorems mentioned above.\[\sqrt{\frac{ 2x^2-2x+1 }{ 2 }} \ge \frac{ 1 }{ x+\frac{ 1 }{ x } }\]\[\sqrt{\frac{ 2\left( x^2-x+\frac{ 1 }{ 2 } \right) }{ 2 }} \ge \frac{ x }{ x^2+1 }\]\[x^2-x+\frac{ 1 }{ 2 } \ge \frac{ x^2 }{ x^4+2x^2+1 }\]\[x^6-x^5+\frac{ 5 }{ 2 }x^4-2x^3+x^2-x+\frac{ 1 }{ 2 } \ge 0\]\[\left( x^6+\frac{ 5 }{ 2 }x^4+x^2+\frac{ 1 }{ 2 } \right) - \left( x^5+2x^3+x \right) \ge 0\]\[x=0 \rightarrow \frac{ 1 }{ 2 } \ge 0\]\[x=1 \rightarrow 0 \ge 0\]

OpenStudy (anonymous):

Excuse me. Last line above \[x=1 \rightarrow 1 \ge 0\]should read

OpenStudy (anonymous):

@ganeshie8 isn't the minimum of the LHS \(\sqrt\frac{1}{2}\) ? If so, how does that affect the max/min proposition?